# Dynamic Contextual Pricing with Doubly Non-Parametric Random Utility Models

### v2

@@ 3.2 Regret Analysis of Decision Making
When $F(\cdot)$ is comparatively more difficult to learn, i.e. less smooth such that $m<1+\min(\frac{12}{d},9)$, the overall regret is dominated by $\widetilde{{\cal O}}(T^{\frac{2m+1}{4m-1}})$. When $\mathring{\mu}(\boldsymbol{x})$ is comparatively more difficult to learn, i.e. higher-dimensional such that $m\geq 1+\min(\frac{12}{d},9)$, the the overall regret is dominated by $\widetilde{{\cal O}}(T^{\frac{d+8}{d+16}\lor\frac{7}{13}})$ where the term $\frac{7}{13}$ corresponds to the setting of $d=1$.
@@
If we only consider the explore-and-exploit algorithms and the policy set $\Pi:=\{\pi:\pi(p_{t})=\widehat{\mu}_{k}(\boldsymbol{x}_{t})+\widehat{\phi}_{k}^{-1}(-\widehat{\mu}_{k}(\boldsymbol{x}_{t}))\}$ with $\widehat{\phi}_{k}(z)=z-\frac{1-\widehat{F}_{k}(z)}{\widehat{F}_{k}^{(1)}(z)}$, where $\widehat{\mu}_{k}$, $\widehat{F}_{k}$, and $\widehat{F}^{(1)}_{k}$ are estimators for the mean utility function, noise distribution, and derivative of the noise distribution in the $k$-th episode, respectively, then the optimality of the regret upper bound reduces to the optimality of these estimators. Based on Stone (1982), our nonparametric estimators DNN or TDNN $\widehat{\mu}_{k}$ , Nadaraya-Watson kernel estimator $\widehat{F}$ and its derivative $\widehat{F}^{(1)}$, all achieve optimal rate of convergence. Therefore, the proposed algorithm using those estimators is minimax optimal when we only focus on the policy set $\Pi$.
@@ 4.1 Synthetic Data
Figure 1: (a): The x-axis is T while the y-axis is $\operatorname{Reg}(\pi,T)$. (b): The x-axis is $\log(T)-\log(1500)$, while the y-axis is $\tilde{\text{reg}}(T)$ defined above. For (a) and (b), the solid lines represent the mean $\operatorname{Reg}(\pi,T)$ and $\tilde{\text{reg}}(T)$, respectively, over 20 independent runs. The light color areas around those solid lines depict the standard error of the estimations. The blue, green, and red lines represent linear regression, TDNN, and DNN estimations in the exploration phase in Algorithm 1, respectively. The orange line represents applying ’RMLP-2’. The dashed lines represent the benchmark whose slopes are equal to $\frac{2m+1}{4m-1}$ with $m=2$.
@@
From Figure 1, we conclude that under the specified setting, the rates of the empirical cumulative regrets produced by Algorithm 1 (as shown by the solid green and red lines) do not exceed their theoretical counterparts given in Theorems 8(as shown by the dashed lines). At majority of time points, the growth rates of the empirical regrets are very close to those of the theoretical line. This demonstrates the tightness of our theoretical results. We also see that the regrets we achieved are much smaller than those two approaches. As for the comparison with linear regression, TDNN and DNN estimations are robust to the mis-specification of the mean utility function, and in comparison with ’RMLP-2’, our method is robust to the mis-specification of the noise distribution since our algorithm can adapt to all noise distributions in the non-parametric class. From Figure 2, we conclude that under the same setting, Algorithm 1 incorporating TDNN estimation surpasses all other methods in performance, since as the episode length increases, the TDNN method achieves a lower regret during the exploitation phase with more accurate estimation of the underline mean utility function. In contrast, the averaged regret curves of exploitation phase for the two benchmark approaches exhibit minimal decay as the time horizon $T$ increases. Such phenomenon arises from model misspecification and the biased learning of the mean utility function by the two benchmark approaches.
@@ 4.2 Real Data Analysis
Note that one is not able to obtain online responses to any algorithms. Thus, we follow the calibration idea proposed in (Ban and Keskin, 2021; Luo et al., 2024; Wang et al., 2023) to first learn a binary choice model using the entire dataset and then leverage it as the ground truth to conduct numerical experiments. To make calibration model more stable, we process all covariates by standardization and take log of the price variable. One can also remove the extreme values that are outside the 0.01 and 0.99 percentiles to mitigate the influence of outliers. For learning the decision model, we typically fit a generalized logistic model with a quadratic form as the logits; meanwhile, instead of using a conventional logistic function, we adopt a specified sigmoid function with bounded support.
Specifically, denote the contextual covariate as $\boldsymbol{x}_{t}\in\mathbb{R}^{4},\;t=1,2,\dots$ and the quadratic form $\mathring{\mu}(\boldsymbol{x}_{t})=\boldsymbol{\beta}^{\top}\boldsymbol{x}_{t}^{2},\;\boldsymbol{\beta}\in\mathbb{R}^{4}$ for the mean utility function. Then the noise $\{\varepsilon_{t}\}$ are set to independently follow the distribution with probability density function $f(z)\propto(1-(\frac{z}{3})^{2})^{4}\cdot\mathbbm{1}(\left\lvert z\right\rvert\leq 3).$ Furthermore, with the belief that the coefficients of the mean utility function can be varied across the data while the noise shares the same distribution. We employ a two-stage approach learning calibration model. Initially, the covariates $\{\boldsymbol{x}_{t}\}_{t\geq 1}$ are divided into 20 clusters based on similarities, utilizing the K-means clustering technique. Here we denote $\boldsymbol{x}_{t}^{(k)}$ if the covariate $\boldsymbol{x}_{t}$ belongs to the $k$-th cluster. Then, we apply generalized logistic regression to each data cluster, fitting the coefficients of the quadratic form for each group while maintaining the same noise distribution. Figure 3 demonstrates the alignment between observed data and the theoretical model over the noise $\varepsilon_{t}$, which means the calibration model is well specified.
@@
Given these key components, the remaining experiments remain almost the same as discussed in the simulation part. Here the calibration model adopts the form $\mathring{\mu}(\boldsymbol{x}_{t})=\mathring{\mu}(\boldsymbol{x}_{t}^{(k)})=\boldsymbol{\beta}_{k}^{\top}\big{(}\boldsymbol{x}_{t}^{(k)}\big{)}^{2},\;k=1,\dots,20$, where $\boldsymbol{\beta}_{k}$ is the learned coefficients when the contextual covariate $\boldsymbol{x}_{t}$ belongs to the $k$-th cluster. Since in reality we don’t know the true smoothness order of noise distribution, in each episode, we set the exploration phase with length $|{\cal T}_{k,exp}|=\lfloor(dn_{k})^{\frac{d+4}{d+8}}\rfloor$. We set $B=6,n_{0}=200$ and conduct Algorithm 1. Recall from Theorem 8 that the regret $\operatorname{Reg}(\pi,T)\lesssim[(\log T)^{2}+d\log T]T^{\frac{d+4}{d+8}}.$ Hence, we plot $\widetilde{\textrm{reg}}(T)$ against $\log(T)-\log(1500)$ in Figure 4(b), where $\widetilde{\textrm{reg}}(T):=\log(\operatorname{Reg}(T))-(2\log\log T+\log 4\log T)-(\log(\operatorname{Reg}(T=1500))-2\log\log 1500-\log 4\log 1500)$.
Figure 4: Comparison between our methods and linear model based policy and ‘RMLP-2’ based on real data application. (a): The x-axis is T while the y-axis is $\operatorname{Reg}(\pi,T)$. (b): The x-axis is $\log(T)-\log(1500)$, while the y-axis is $\tilde{\text{reg}}(T)$ defined above. The dashed lines has slope equal to $\frac{d+4}{d+8}$ with $d=4$.

### v3

@@ 3.2 Regret Analysis of Decision Making
When $F(\cdot)$ is comparatively more difficult to learn, i.e. less smooth in the sense that $m<1+\min(\frac{12}{d},9)$, the overall regret is dominated by $\widetilde{{\cal O}}(T^{\frac{2m+1}{4m-1}})$. When $\mathring{\mu}(\boldsymbol{x})$ is comparatively more difficult to learn, i.e. higher-dimensional such that $m\geq 1+\min(\frac{12}{d},9)$, the the overall regret is dominated by $\widetilde{{\cal O}}(T^{\frac{d+8}{d+16}\lor\frac{7}{13}})$ where the term $\frac{7}{13}$ corresponds to the setting of $d=1$.
Remark 3.
The structure of the random utility model leads to (16) and (17), facilitating an advanced analysis of regret, which demonstrates a high-order dependence on the square of the difference between the optimal and estimated prices. To elaborate, consider the revenue function $\operatorname{rev}_{t}(p)$ defined in (3) along with its first and second derivatives, $\operatorname{rev}_{t}^{(1)}(\cdot)$ and $\operatorname{rev}_{t}^{(2)}(\cdot)$ respectively. As $p_{t}^{*}$ is the price that maximizes revenue, it holds that $\operatorname{rev}_{t}^{(1)}(p_{t}^{*})=0$. A Taylor expansion around $p_{t}^{*}$ yields
$\operatorname{rev}_{t}(p_{t})=\operatorname{rev}_{t}(p_{t}^{*})+\frac{1}{2}\operatorname{rev}_{t}^{(2)}(p)(p_{t}-p_{t}^{*})^{2},$
for some $p$ between $p_{t}$ and $p_{t}^{*}$. Additionally, according to (17), the $(p_{t}-p_{t}^{*})^{2}$ is directly linked to the square of the estimation errors in $F$ and $\mathring{\mu}$.
This nuanced relationship results in improved regret bounds under various scenarios, surpassing those reported in Chen and Gallego (2021) and Bu et al. (2022). Unlike the linear dependence on the difference between optimal and estimated prices observed in the discretization-based methods used in these studies, our model’s regret depends quadratically on these differences, leading to potentially tighter regret bounds under comparable conditions.
@@
When we consider the policy set $\Pi:=\{\pi:\pi(p_{t})=\widehat{\mu}_{k}(\boldsymbol{x}_{t})+\widehat{\phi}_{k}^{-1}(-\widehat{\mu}_{k}(\boldsymbol{x}_{t}))\}$ with $\widehat{\phi}_{k}(z)=z-\frac{1-\widehat{F}_{k}(z)}{\widehat{F}_{k}^{(1)}(z)}$, where $\widehat{\mu}_{k}$, $\widehat{F}_{k}$, and $\widehat{F}^{(1)}_{k}$ are estimators for the mean utility function, noise distribution, and derivative of the noise distribution in the $k$-th episode, respectively, then the optimality of the regret upper bound reduces to the optimality of these estimators. Based on Stone (1982), our nonparametric estimators DNN or TDNN $\widehat{\mu}_{k}$ , Nadaraya-Watson kernel estimator $\widehat{F}$ and its derivative $\widehat{F}^{(1)}$, all achieve optimal rate of convergence. Therefore, the proposed algorithm using those estimators is minimax optimal when we only focus on the policy set $\Pi$.
@@ 4.1 Synthetic Data
Figure 1: (a): The x-axis is T while the y-axis is $\operatorname{Reg}(\pi,T)$. (b): The x-axis is $\log(T)-\log(1500)$, while the y-axis is $\tilde{\text{reg}}(T)$ defined above. For (a) and (b), the solid lines represent the mean $\operatorname{Reg}(\pi,T)$ and $\tilde{\text{reg}}(T)$, respectively, over 20 independent runs. The light color areas around those solid lines depict the standard error of the estimations. The blue, green, and red lines represent linear regression, TDNN, and DNN estimations in the exploration phase in Algorithm 1, respectively. And the orange line represents applying ’RMLP-2’. The dashed lines represent the benchmark whose slopes are equal to $\frac{2m+1}{4m-1}$ with $m=2$.
@@
From Figure 1, we conclude that under the specified setting, the rates of the empirical cumulative regrets produced by Algorithm 1 (as shown by the solid green and red lines) do not exceed their theoretical counterparts given in Theorems 8(as shown by the dashed lines). In many cases, the growth rates of the empirical regrets are very close to those of the theoretical line. This demonstrates the tightness of our theoretical results. We also see that the regrets we achieved are much smaller than those two approaches. As for the comparison with linear regression, TDNN and DNN estimations are robust to the mis-specification of the mean utility function, and in comparison with ’RMLP-2’, our method is robust to the mis-specification of the noise distribution since our algorithm can adapt to all noise distributions in the non-parametric class. From Figure 2, we conclude that under the same setting, Algorithm 1 incorporating TDNN estimation surpasses all other methods in performance, since as the episode length increases, the TDNN method accelerates the convergence of regret during the exploitation phase with more accurate estimation of the underline mean utility function and consequently, more accurately approaching towards noise distribution. In contrast, the decay in the averaged regret of exploitation phase for the two benchmark approaches nearly goes to zero. This occurs because the increasing size of sample points during the exploration phase does not enhance the estimation accuracy of the mean utility function due to mis-specification.
@@ 4.2 Real Data Analysis
Note that one is not able to obtain online responses to any algorithms. Thus, we follow the calibration idea proposed in (Ban and Keskin, 2021; Luo et al., 2024; Wang et al., 2023) to first learn a binary choice model using the entire dataset and then leverage it as the ground truth to conduct numerical experiments. To make calibration model more stable, we process all covariates by standardization and take log of the price variable. One can also remove the extreme values that are outside the 0.01 and 0.99 percentiles to mitigate the influence of outliers. For learning the decision model, we typically fit a generalized logistic model with a quadratic form as the logits; meanwhile, instead of using a conventional logistic function, we adopt a specified sigmoid function with bounded support. Specifically, denote the contextual covariate as $\boldsymbol{x}_{t}\in\mathbb{R}^{4},\;t=1,2,\dots$ and the quadratic form $\mathring{\mu}(\boldsymbol{x}_{t})=\boldsymbol{\beta}^{\top}\boldsymbol{x}_{t}^{2},\;\boldsymbol{\beta}\in\mathbb{R}^{4}$ for the mean utility function. Then the noise $\{\varepsilon_{t}\}$ are set to independently follow the distribution with probability density function $f(z)\propto(1-(\frac{z}{3})^{2})^{4}\cdot\mathbbm{1}(\left\lvert z\right\rvert\leq 3).$ Furthermore, with the belief that the coefficients of the mean utility function can be varied across the data while the noise shares the same distribution. We employ a two-stage approach learning calibration model. Initially, the covariates $\{\boldsymbol{x}_{t}\}_{t\geq 1}$ are divided into 20 clusters based on similarities, utilizing the K-means clustering technique. Here we denote $\boldsymbol{x}_{t}^{(k)}$ if the covariate $\boldsymbol{x}_{t}$ belongs to the $k$-th cluster. Then, we apply generalized logistic regression to each data cluster, fitting the coefficients of the quadratic form for each group while maintaining the same noise distribution. Figure 3 demonstrates the alignment between observed data and the theoretical model over the noise $\varepsilon_{t}$, which means the calibration model is well specified.
@@
Given these key components, the remaining experiments remain almost the same as discussed in the simulation part. Here we use the calibration model, setting $\mathring{\mu}(\boldsymbol{x}_{t})$ with parameter $\boldsymbol{\beta}_{k},k=1,\dots,20$, and the noise distribution $f(v)$ given above and sample $\boldsymbol{x}_{t}$ from those four features. Since in reality we don’t know the true smoothness order of noise distribution, in each episode, we set the exploration phase with length $|{\cal T}_{k,exp}|=\lfloor(dn_{k})^{\frac{d+4}{d+8}}\rfloor$. We set $B=6,n_{0}=200$ and conduct Algorithm 1. Recall from Theorem 8 that the regret $\operatorname{Reg}(\pi,T)\lesssim[(\log T)^{2}+d\log T]T^{\frac{d+4}{d+8}}.$ Hence, we plot $\widetilde{\textrm{reg}}(T)$ against $\log(T)-\log(1500)$ in Figure 4(b), where $\widetilde{\textrm{reg}}(T):=\log(\operatorname{Reg}(T))-(2\log\log T+\log 4\log T)-(\log(\operatorname{Reg}(T=1500))-2\log\log 1500-\log 4\log 1500)$.
Figure 4: Comparison between our methods and linear model based policy and ‘RMLP-2’ based on real data application. (a): The x-axis is T while the y-axis is $\operatorname{Reg}(\pi,T)$. (b): The x-axis is $\log(T)-\log(1500)$, while the y-axis is $\tilde{\text{reg}}(T)$ defined above. The dashed lines has slope equal to $\frac{d+4}{d+8}$ with $d=4$. Here we only run algorithm one time.